OpenStudy (goformit100):

Which of the following is not a way to represent the solution of the inequality 2(x + 1) greater than or equal to 20? x 9 9 x A number line with a closed circle on 9 and shading to the left. A number line with a closed circle on 9 and shading to the right.

OpenStudy (goformit100):

@texaschic101 HELP

OpenStudy (goformit100):

@Nurali HELP ME MADAm.

OpenStudy (texaschic101):

2(x + 1) >= 20 2x + 2 >= 20 2x >= 20 - 2 2x >= 18 x >= 9 I can't see the variables in the first two answer choices, however, the third answer choice is incorrect.